OpenStudy (anonymous):

A student drops a rock from a bridge to the water 12 m below. With what speed does the rock strike the water? Answer in units of m/s

OpenStudy (anonymous):

v^2 = u^2 + 2as a in this case is g = 9.81 or 10 depending what accuracy you want, and you know s, the distance. U is the starting velocity and I guess that this is zero.

OpenStudy (anonymous):

Use the equation\[X(12m)=X(0)+V(original).time+gravity.{time^{2}}\]after cancelling, it becomes\[X(12m)=gravity.{time^{2}}\]solved for time\[Time(s)=\sqrt{X(12m)/gravity}\]then use the equation \[V(final)=gravity.\sqrt{X(12m)/gravity}\]which becomes\[V(final)=\sqrt{X(12m)gravity}\]gravity constant is roughly rounded to 10meters/second square and all your units are already in metric system.

OpenStudy (anonymous):

For those of you struggling with questions about things moving under constant accelerations - like gravity - the equations you need are very easy - but only if you do a few examples yourself. For constant accelerations v=u+a.t s=u.t + 1/2 * a.t^2, and v^2 = u^2 + 2.a.s where u= initial velocity v = final velocity s = distance a =acceleration and t = time